OpenStudy (anonymous):

Find the area of a circle whose area equals the half of the sum of the areas of two circles with radii 2 and 3. the way this is written is confusing me can someone please explain it to me

OpenStudy (anonymous):

i want to learn and me just getting the answer off the internet wont help me at all

OpenStudy (anonymous):

thank you for the quick response

OpenStudy (anonymous):

I put it in more "mathematical "notation, :) "A" will be the area you looking for. The circle are formula is: Pi*r^2. So..... A=1/2(Pi*2^2+Pi*3^2) Ok?

OpenStudy (anonymous):

ok give me ona momenta to solve

OpenStudy (anonymous):

radii is the radius right :/

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

12.5 pie?

jimthompson5910 (jim_thompson5910):

The area of the first circle (with a radius of 2) is A = pi*r^2 A = pi*2^2 A = pi*4 A = 4pi The area of the second circle (with a radius of 3) is A = pi*r^2 A = pi*3^2 A = pi*9 A = 9pi ------------------------------------------------------- The sum of the two areas is just the result of adding the two areas up. So Sum of two areas = (Area 1) + (Area 2) Sum of two areas = 4pi + 9pi Sum of two areas = 13pi Now take half of this sum to get 13pi/2 = 6.5pi

OpenStudy (anonymous):

A=1/2(Pi*2^2+Pi*3^2) = 1/2(13*Pi)

OpenStudy (anonymous):

ohh! ok i see so i used 4 not 2 thank you both!!

jimthompson5910 (jim_thompson5910):

yw